OpenStudy (bloomlocke367):

Eli has 7 black pens and 5 blue pens in his desk drawer. He also has 3 yellow highlighters, 2 green highlighters, and 5 pink highlighters in his pencil case. If he chooses one pen and one highlighter without looking, what is the probability that he will get a black pen and he will not get a green highlighter?

OpenStudy (bloomlocke367):

@rational

OpenStudy (rational):

consider the events separately

OpenStudy (rational):

` 7 black pens and 5 blue pens` what is the probability that he will get a black pen ?

OpenStudy (bloomlocke367):

7/12?..

OpenStudy (anonymous):

yep

OpenStudy (rational):

Yes. ` 3 yellow highlighters, 2 green highlighters, and 5 pink highlighters ` whats the probability that he will get something other than green ?

OpenStudy (anonymous):

2/10 for highlighters

OpenStudy (rational):

careful, we don't want green

OpenStudy (bloomlocke367):

it's 8/10

OpenStudy (bloomlocke367):

so 4/5

OpenStudy (rational):

Correct! multiply both

OpenStudy (anonymous):

I mean 8 out 10

OpenStudy (bloomlocke367):

7/15?

OpenStudy (bloomlocke367):

that's it?

OpenStudy (rational):

probability that `he will get a black pen` and `he will not get a green highlighter` = \[\dfrac{7}{12}\times \dfrac{4}{5} \]

OpenStudy (rational):

yeah that simplifies to \[\large \dfrac{7}{15}\]

OpenStudy (rational):

thats it, good job !

OpenStudy (bloomlocke367):

okay, thanks!